OpenStudy (anonymous):

Given the equation Square root of 8x plus 1 = 5, solve for x and identify if it is an extraneous solution. A)x = One fourth, solution is not extraneous B)x = One fourth, solution is extraneous C)x = 3, solution is not extraneous D)x = 3, solution is extraneous

terenzreignz (terenzreignz):

\[\Large \sqrt{8x+1}=5\] is this it?

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

Well then, how do you go about solving for x?

OpenStudy (anonymous):

would u square both sides to get the 8x out the square root?

terenzreignz (terenzreignz):

Squaring both sides seems logical...what do you get? :)

OpenStudy (anonymous):

(8x+1)^2 = 5^2 64x+1= 25?

OpenStudy (anonymous):

I don't know if I did that right

terenzreignz (terenzreignz):

Oh, okay, two things, first, and this is very important (though actually irrelevant to your problem) \[\Large (8x+1)^2 \color{red}\ne 64x + 1\] Do you remember FOIL? \[\Large (8x+1)^2 = (8x+1)(8x+1)=64x^2 + 16x+1\] But anyway, again, that's not relevant to this problem... do recall that squaring and the *radical sign* cancel each other out... like so: \[\Large (\sqrt a)^2=a\]

OpenStudy (anonymous):

kinda

terenzreignz (terenzreignz):

kinda? When you put it that way, it isn't enough :P \[\Large (\sqrt a)^2 = a\]\[\Large \left(\sqrt{8x+1}\right)^2 = \color{red}?\]

OpenStudy (anonymous):

wouldnt it be 64x^2 + 16x +1, but what do I do after that

terenzreignz (terenzreignz):

No, it won't :P Look again at my previous post... I said that thing involving 64x^2 + 16x + 1 has nothing to do with your current problem :)

OpenStudy (anonymous):

so since they cancel itself outitd be just eithe just x or 8x+1

terenzreignz (terenzreignz):

Which is it? ^_^ (Sorry for egging you, this is necessary for self-reliance...)

OpenStudy (anonymous):

8x+1 :D

OpenStudy (tkhunny):

Keep in mind that for this to be the correct result, we MUST know that \(8x + 1 \ge 0\;or\; x \ge -1/8\). This is not an optional thing to know.

terenzreignz (terenzreignz):

So after you do square both sides, you get...?

OpenStudy (anonymous):

8x+1=25????

terenzreignz (terenzreignz):

That's good. Now solve for x ^_^

OpenStudy (anonymous):

x=3

terenzreignz (terenzreignz):

Now you have a tentative value for x... is it extraneous, though? :)

terenzreignz (terenzreignz):

It mustn't result in a radical that I like to call... 'illegal'. Plug it into the radicand 8x+1 and make sure you get something that is not a negative number.... so, DO you get a negative number when you plug in x = 3 into the radicand?

OpenStudy (anonymous):

no... and extraneous solution is a invalid solution right?

terenzreignz (terenzreignz):

Yup... so your final answer is choice.... which choice? :)

OpenStudy (anonymous):

C). x=3 not extraneous

OpenStudy (anonymous):

i got it right on the test THANK YOU!

terenzreignz (terenzreignz):

Well... that wasn't too hard, was it? ^_^ Good job.